Microsoft Stock has a beta of 1.20.

(1) If the market return increased by 15%, what impact would this change be expected to have on Microsoft stock’s return?

(2) If the market return decreased by 8%, what impact would this change be expected to have on Microsoft stock’s return?

(3) Would Microsoft stock be considered more or less risky than the market? Explain
Answer #1

1) The Expected change in Microsoft return =Beta*market return =1.20*15% =18%

2) The Expected change in Microsoft return =-Beta*market return =1.20*8% =-9.6%

3) Yes it is risky because beta is greater than 1. So its risk is greater than market risk. In case of recession its return is less.